Given two parallel metal plates of dimensions, 20cm by 40cm and each carrying an electric charge of 0.2micro coulomb's. Calculate the electric flux density on each plate and the electric field strength when the plates are placed 5mm apart and the voltage across them is 250volts.
1
Expert's answer
2021-09-08T17:59:18-0400

Electric flux density

"D=\frac{Q}{4\pi r^2}"

"D=\frac{Q}{A}"


"D=\frac{0.2\times10^{-6}}{20\times40\times10^{-4}}=0.25\times10^{-5}c/m^2"

Electric field

"E=\frac{V}{l}=\frac{250}{5\times10^{-3}}=5\times10^4V/m"